OpenStudy (anonymous):

Solve log_2 (3 − x) + log_2 (−x) = 2.

OpenStudy (anonymous):

log2(3x-x^2)=2 2^2 = 3x-x^2 0=-x^2+3x-4 x=4,-1 check for validity,

OpenStudy (anonymous):

log_2(-3x+x²)=2 log_2(-3x+x²)=lg_2(4) x²-3x-4=0 (x-4)(x+1)=0 x=4 or -1

OpenStudy (anonymous):

Use the change of base formula to find log3 ()

OpenStudy (anonymous):

Use the change of base formula to find log_3 (pi)

OpenStudy (anonymous):

Use the change of base formula to find log_3 (pi)? different question?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

\[\log_{10} 3\Pi \div \log_{10} 10\]

OpenStudy (anonymous):

or ln3(pi)/ln10

OpenStudy (anonymous):

plug either into your calculator

OpenStudy (anonymous):

the options i have are a) log3/log pi b)log(pi/3) c) log pi/log3 d) log(3/pi) e) log_pi(^3)

OpenStudy (anonymous):

is the question log(base 3) (pi)? if so it would C

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

"C"

OpenStudy (anonymous):

logbA= logA/logb you digg?

OpenStudy (anonymous):

yea lol im doing some review for my math exam i really appreciate the help i have more questions if you dont mind

OpenStudy (anonymous):

im studying myself.. if i can, i will answer your questions